A father is 24 years older than his son. In 8years he will be twice as old as his son. Determine their present ages. Please, Thank you~
Answer by scott8148(6628)

"A father is 24 years older than his son" ___ f = s + 24

"In 8 years he will be twice as old as his son" ___ f + 8 = 2(s + 8)

substituting ___ (s + 24) + 8 = 2s + 16

solve for s, and substitute back to find f